Grade 11General Physics

If velocity of light c, Planks constant h and Gravitational constant G are taken as fundamental quantites . Express mass in terms of dimensions of these physical quantites.

Sourav R Manoj
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
C = L T⁻¹
h = Joules second = energy / frequency = M L² T⁻¹
G = Newtons meter² / kg²  = M⁻¹ L³ T⁻² 
M = c^k  h^m  G^n  = M^(m-n)  L^(k+2m+3n)  T^(-k-m -2n)
    =>  k+2m+3n = 0,   m - n = 1,    and      - k - m - 2 n = 0
    =>  2m +3n = m + 2n,   m = -n    and   so  n = -1/2,   m = 1/2   and  k = 1/2
                   M = √c  √h / √G
 
 
Regards
Arun (askIITians forum expert)